OpenStudy (anonymous):

Amie had an operation on her knee. To help it get strong again she walks around this rectangular path in the park every morning. If each step she takes is 24 inches long, how many steps does she take on her walk?

OpenStudy (anonymous):



OpenStudy (anonymous):

, so would it be 325, 650, 900, or 1,300 steps?

OpenStudy (anonymous):

\[12\]inches make a foot, find the perimeter

OpenStudy (anonymous):

ok well her steps one 1 equals 24 ft, still trying to figure this out, unsure of exactly how I'm supposed to find the perimeter

OpenStudy (anonymous):

@KendrickLamar2014

OpenStudy (anonymous):

@mathmath333

OpenStudy (anonymous):

Her step equals 2 feet not 24. Find the perimeter by adding all the values in your rectangle, then divide by what how many feet her step equals, you should find a smaller number than the perimeter since her step is longer and therefore she uses less steps to walk the entire path

OpenStudy (anonymous):

1300 I think

OpenStudy (mathmath333):

for i complete round she has to walk perimeter of the park p=450*2+200*2=1300

OpenStudy (mathmath333):

in foots

OpenStudy (mathmath333):

p=1300*12 in inches her one step =24 inches therefore total steps=1300*12/24=650

OpenStudy (anonymous):

Thank You!